# Charm physics with overlap fermions on 2+1-flavor domain wall fermion configurations

### v1

@@ Abstract
Decay constants of pseudoscalar mesons $D$, $D_{s}$, $\eta_{c}$ and vector mesons $D^{*}$, $D_{s}^{*}$, $J/\psi$ are determined from $N_{f}=2+1$ lattice QCD at a lattice spacing $a\sim 0.08$ fm. For vector mesons, the decay constants defined by tensor currents are given in the ${\overline{\rm MS}}$ scheme at $2$ GeV. The calculation is performed on domain wall fermion configurations generated by the RBC-UKQCD Collaborations and the overlap fermion action is used for the valence quarks. Comparing the current results with our previous ones at a coarser lattice spacing $a\sim 0.11$ fm gives us a better understanding of the discretization error. We obtain $f_{D_{s}^{*}}^{T}({\overline{\rm MS}},\mbox{ 2 GeV})/f_{D_{s}^{*}}=0.907(20)$ with a better precision than our previous result. Combining our $f_{D_{s}^{*}}=277(11)$ MeV with the total width of $D_{s}^{*}$ determined in a recent work gives a branching fraction $4.26(52)\times 10^{-5}$ for $D_{s}^{*}$ leptonic decay.
1Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China
2School of Physical Sciences, University of Chinese Academy of Sciences, Beijing 101408, China
3Department of Physics and Astronomy, University of Kentucky, Lexington, Kentucky 40506, USA
4Nuclear Science Division, Lawrence Berkeley National Laboratory, Berkeley, California 94720, USA
5Center for High Energy Physics, Peking University, Beijing 100871, China
@@ 1 Introduction
Charm physics provides a rich phenomenology and offers us a valuable platform for precisely testing the Standard Model and understanding low-energy Quantum Chromodynamics (QCD). The decay constants of charmed pseudoscalar mesons combined with experimental data of the relevant leptonic decays can be used to extract the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements $V_{cd}$ and $V_{cs}$ (see [1]), or serve as probes for new particles of new physics such as charged Higgs bosons (see [2]). However, perturbative theory and heavy quark expansion are rather precarious in charm sector, and our most reliable theoretical tool is lattice QCD. An overview of the decay constants of charmed pseudoscalar mesons can be found in Ref. [3] for lattice QCD calculations before 2022. Two recent $2+1$-flavor calculations of these decay constants appeared in Refs. [4, 5].
@@
For the pseudoscalar meson $D$ or $D_{s}$, its pure leptonic decay $D_{(s)}\to\ell\nu_{\ell}$, of which the decay width is proportional to $f^{2}_{P}|V_{cq}|^{2}$ ($q=d\mbox{ or }s$ with $f_{P}$ the decay constant of $D\mbox{ or }D_{s}$, respectively), provides a clean channel to determine the CKM matrix elements. Moreover, the semileptonic processes $D_{(s)}\to h\ell\nu_{\ell}$, with $h$ representing a pion or kaon, are induced by the vector current $\bar{q}\gamma^{\mu}c$, while the pure leptonic processes are induced by the axial-vector current $\bar{q}\gamma^{\mu}\gamma_{5}c$. Comparing the CKM matrix elements extracted from these two kinds of processes allows us to test the $V-A$ structure of Weak interaction. The vector mesons $D_{(s)}^{*}$ almost 100% decay respectively to the corresponding pseudoscalar charmed mesons, with relatively small leptonic decay branching ratios. The first measurement of $f_{D_{s}^{*}}$ from the leptonic decay $D_{s}^{*}\to e^{+}\nu_{e}$ is recently reported by BESIII [6]. In the phenomenological analysis of the nonleptonic weak decays of charmed vector mesons, the decay constants denoted by $f_{V}$ are necessary inputs [7, 8]. In each of the $b\to c$ induced semileptonic or nonleptonic decays of bottom mesons, the decay constant $f_{V}$ of the charmed meson in the final state, as well as the decay constant $f_{V}^{T}$ defined by the tensor current, appears as nonperturbative inputs [9, 10]. Furthermore, the ratio of the decay constants of vector mesons and pseudoscalar mesons $f_{V}/f_{P}$ approaches to one in the heavy quark limit, which can be used to study the breaking of heavy quark symmetry. Currently, unquenched lattice QCD calculations of decay constants of charmed vector mesons are relatively sparse [11, 12, 13, 14, 15].
@@
For the charmonium system, the leptonic decay constant of $J/\psi$ can be directly determined experimentally [16]. In this work, we calculated the ratios of decay constants $f_{J/\psi}/f_{\eta_{c}}$ and $f^{T}_{J/\psi}/f_{J/\psi}$, where the former can be directly used in the amplitude analysis of $B$ meson decays into charmonium states, and the latter will appear in $b\to c$ decays induced by specific new physics operators. By combining the decay width $\Gamma(J/\psi\to e^{+}e^{-})$ measured experimentally with the ratio $f_{J/\psi}/f_{\eta_{c}}$, we can determine the decay constant $f_{\eta_{c}}$, which is used for the calculation of decay amplitudes $\Gamma(\eta_{c}\to\gamma^{*}\gamma)$ and form factors appearing in $\gamma^{*}\gamma^{*}\to\eta_{c}$ processes [17, 18, 19]. Theoretically, the decay constants $f_{J/\psi}$ and $f_{\eta_{c}}$ were calculated in quenched lattice QCD in [20]. Two flavor lattice QCD calculations of $f_{J/\psi}$ and $f_{\eta_{c}}$ are given in [21] and [22]. The HPQCD collaboration obtains $f_{J/\psi}$ and/or $f_{\eta_{c}}$ in $2+1$-flavor [23, 24], and $2+1+1$-flavor simulations [25]. The latter work also considers the quenched Quantum Electrodynamics (QED) effects of charm quarks.
@@ 2 Simulation details
The physical mass points of light, strange and charm quarks are fixed by using the meson masses $m_{\pi}^{\rm expt}$, $m_{K}^{\rm expt}$ and $m_{J/\psi}^{\rm expt}$ measured experimentally. Note that at the physical pion mass, $D^{*}$ decays into the $P$-wave $D\pi$ state through the strong interaction. In Table 3, we have listed some masses of $D^{(*)}$, with different valence pion masses on our lattice. It is seen that, at each valence pion mass, the mass of the vector meson $D^{*}$ is below the $D\pi$ threshold. Therefore, in this study $D^{*}$ is stable on our lattice. Similarly, $D_{s}^{*}$ is also a stable particle.
@@ 3.1 Meson masses
Figure 2: Left: fitted $aM_{D_{s}}$ as a function of $t_{\text{min}}$ with fixed $t_{\text{max}}=32$. Middle: effective masses (blue points) from the two-point function, compared with $aM_{D_{s}}$ from the fit (orange band) with fitting range $15\leq t\leq 32$. Right: fitted $aM_{D_{s}}$ as a function of the number of states from the multi-state fit.
The right panel of Fig. 2 shows $aM_{D_{s}}$ from the multi-state fits as a function of the number of states $n$ in the fitting function [37]. After $n$ exceeds $4$, the fitting results stabilize and agree with that from the one-state fit.
@@
We also obtained the masses of pion and kaon. Then we use the experimentally measured $m_{\pi}^{\rm expt}$, $(m_{ss}^{\rm phys})^{2}\equiv 2(m_{K}^{\rm expt})^{2}-(m_{\pi}^{\rm expt})^{2}$, and $m_{J/\psi}^{\rm expt}$ as inputs to determine the physical meson masses and decay constants. From Tab. 5, it can be seen that $(am_{\pi})^{2}$ shows a linear dependence on the light quark mass, while $(am_{ss})^{2}$ is independent of it but shows a linear dependence on the strange quark mass, which are consistent with the expectations of chiral perturbation theory.
@@
After obtaining the meson masses and decay constants at different valence quark masses on ensemble f004, we used linear interpolation (extrapolation) to obtain the results at the physical mass point. The Particle Data Group [1] gives
$m_{\pi}^{\rm expt}=139.57\ \text{MeV},\quad m_{K}^{\rm expt}=493.67\ \text{MeV},\quad m_{J/\psi}^{\rm expt}=3.0969\ \text{GeV}.$
@@
$(am_{\pi})^{2}_{\rm phys}=0.00343(3)\,,\ (am_{ss})^{2}_{\rm phys}=0.0824(7)\,,\ (am_{J/\psi})_{\rm phys}=1.2996(49)\,.$
@@
$M_{D}^{\text{phys}}=1.866(14)(7)\,\text{GeV},\quad M_{D^{*}}^{\text{phys}}=2.008(24)(8)\,\text{GeV},$ (17)
$M_{D_{s}}^{\text{phys}}=1.9659(46)(75)\,\text{GeV},\quad M_{D_{s}^{*}}^{\text{phys}}=2.1127(71)(80)\,\text{GeV},$
@@ 3.2 Decay constants
$f_{D}=215.4(9.0)(0.9),\quad f_{D^{*}}=224.3(16.0)(0.9),\quad f_{D^{*}}^{T}=191.5(12.5)(1.9)(0.8),$ (21)
$\,f_{D_{s}}=255.4(4.3)(1.0),\quad f_{D_{s}^{*}}=276.7(6.4)(1.1),\quad f_{D_{s}^{*}}^{T}=251.8(6.0)(2.4)(1.0).$
Here, for each of the decay constants, the first error includes statistical and extrapolation/interpolation uncertainties, the last error is from the uncertainty of the lattice spacing, and the error in the middle, if there is, is due to the uncertainty in the renormalization constant. It is seen in Eq. (8) that the uncertainty of $Z_{V}(=Z_{A})$ can be regarded as negligible. The uncertainty of $Z_{T}$ is around 1% and is much smaller than the first error.
For chiral fermions as used in this work the discretization effects are proportional to squared lattice spacing $a^{2}$. For $f_{D_{s}^{(*)}}$, our results (at $a^{2}\approx 0.007\,\mathrm{fm}^{2}$) are consistent within $1\sigma$ deviation with the previous results (at $a^{2}\approx 0.013\,\mathrm{fm}^{2}$), respectively, where $f_{D_{s}}=249(5)\,$MeV and $f_{D_{s}^{*}}=274(5)\,$MeV [26] and we have dropped the discretization error assigned to the previous work. For $f_{D^{(*)}}$, our results are also consistent within $1\sigma$ deviation with the previous results, respectively, where $f_{D}=213(2)\,$MeV and $f_{D^{*}}=234(3)\,$MeV. The differences in the central values are approximately from 1% to $4\%$. Therefore, we take an average value $3\%$ as the estimate of the discretization errors in our decay constants and get (in units of MeV),
$f_{D}=215.4(9.0)(6.5),\quad f_{D^{*}}=224.3(16.1)(6.8),\quad f_{D^{*}}^{T}=191.5(12.6)(5.8),$ (22)
$\,f_{D_{s}}=255.4(4.4)(7.7),\quad f_{D_{s}^{*}}=276.7(6.5)(8.3),\quad f_{D_{s}^{*}}^{T}=251.8(6.5)(7.6),$
@@
$f_{J/\psi}/f_{\eta_{c}}=1.059(15),\quad f^{T}_{J/\psi}/f_{J/\psi}=0.961(10).$ (23)
Here the error includes statistical uncertainty and the one in renormalization constants. Combining these values with the experimental measured $f_{J/\psi}^{\rm expt}=407(4)\,$MeV, which is obtained from its pure leptonic decay width $\Gamma(J/\psi\to e^{+}e^{-})$ [25], we can derive $f_{\eta_{c}}=384.3(6.6)\,$MeV and $f^{T}_{J/\psi}=391.1(5.6)\,$MeV. These results are consistent within $1.5\sigma$ deviation with the results $f_{\eta_{c}}=394.7(2.4)\,$MeV [23], $f^{T}_{J/\psi}=392.7(2.7)\,$MeV and $f^{T}_{J/\psi}/f_{J/\psi}=0.9569(52)$ [38] from the HPQCD Collaboration obtained in the continuum limit. The better agreement in the ratio means the discretization effects are shrunk by the cancellation between the numerator and the denominator. Because we do not have results of decay constants of charmonia from the previous work to compare, we do not give discretization errors for those constants and ratios in this work.

### v3

@@ Abstract
Decay constants of pseudoscalar mesons $D$, $D_{s}$, $\eta_{c}$ and vector mesons $D^{*}$, $D_{s}^{*}$, $J/\psi$ are determined from $N_{f}=2+1$ lattice QCD at a lattice spacing $a\sim 0.08$ fm. For vector mesons, the decay constants defined by tensor currents are given in the ${\overline{\rm MS}}$ scheme at $2$ GeV. The calculation is performed on domain wall fermion configurations generated by the RBC-UKQCD Collaborations and the overlap fermion action is used for the valence quarks. Comparing the current results with our previous ones at a coarser lattice spacing $a\sim 0.11$ fm gives us a better understanding of the discretization error. We obtain $f_{D_{s}^{*}}^{T}({\overline{\rm MS}},\mbox{ 2 GeV})/f_{D_{s}^{*}}=0.909(18)$ with a better precision than our previous result. Combining our $f_{D_{s}^{*}}=277(11)$ MeV with the total width of $D_{s}^{*}$ determined in a recent work gives a branching fraction $4.26(52)\times 10^{-5}$ for $D_{s}^{*}$ leptonic decay.
1MOE Frontiers Science Center for Rare Isotopes,
and School of Nuclear Science and Technology, Lanzhou University, Lanzhou 730000, China
2Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China
3School of Physical Sciences, University of Chinese Academy of Sciences, Beijing 101408, China
4Department of Physics and Astronomy, University of Kentucky, Lexington, Kentucky 40506, USA
5Nuclear Science Division, Lawrence Berkeley National Laboratory, Berkeley, California 94720, USA
6Center for High Energy Physics, Peking University, Beijing 100871, China
@@ 1 Introduction
Charm physics provides a rich phenomenology and offers us a valuable platform for precisely testing the Standard Model and understanding low-energy Quantum Chromodynamics (QCD). The decay constants of charmed pseudoscalar mesons combined with experimental data of the relevant leptonic decays can be used to extract the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements $V_{cd}$ and $V_{cs}$ (see [1]), or serve as probes for new particles of new physics such as charged Higgs bosons (see [2]). However, perturbative theory and heavy quark expansion are rather precarious in charm sector, and our most reliable theoretical tool is lattice QCD. An overview of the decay constants of charmed pseudoscalar mesons can be found in Ref. [3] for lattice QCD calculations before 2022 (see the individual results in Refs. [4, 5, 6, 7, 8, 9]). Two recent $2+1$-flavor calculations of these decay constants appeared in Refs. [10, 11].
@@
For the pseudoscalar meson $D$ or $D_{s}$, its pure leptonic decay $D_{(s)}\to\ell\nu_{\ell}$, of which the decay width is proportional to $f^{2}_{P}|V_{cq}|^{2}$ ($q=d\mbox{ or }s$ with $f_{P}$ the decay constant of $D\mbox{ or }D_{s}$, respectively), provides a clean channel to determine the CKM matrix elements. Moreover, the semileptonic processes $D_{(s)}\to h\ell\nu_{\ell}$, with $h$ representing a pion or kaon, are induced by the vector current $\bar{q}\gamma^{\mu}c$, while the pure leptonic processes are induced by the axial-vector current $\bar{q}\gamma^{\mu}\gamma_{5}c$. Comparing the CKM matrix elements extracted from these two kinds of processes allows us to test the $V-A$ structure of Weak interaction. The vector mesons $D_{(s)}^{*}$ almost 100% decay respectively to the corresponding charmed pseudoscalar mesons, with relatively small leptonic decay branching ratios. The first measurement of $f_{D_{s}^{*}}$ from the leptonic decay $D_{s}^{*}\to e^{+}\nu_{e}$ is recently reported by BESIII [12]. In the phenomenological analysis of the nonleptonic weak decays of charmed vector mesons, the decay constants denoted by $f_{V}$ are necessary inputs [13, 14]. In each of the $b\to c$ induced semileptonic or nonleptonic decays of bottom mesons, the decay constant $f_{V}$ of the charmed meson in the final state, as well as the decay constant $f_{V}^{T}$ defined by the tensor current, appears as nonperturbative inputs [15, 16]. Furthermore, the ratio of the decay constants of vector mesons and pseudoscalar mesons $f_{V}/f_{P}$ approaches to 1 in the heavy quark limit, which can be used to study the breaking of heavy quark symmetry. Currently, unquenched lattice QCD calculations of decay constants of charmed vector mesons are relatively sparse [17, 18, 19, 20, 21]. Results from sum rules can be found in, for example, Refs. [22, 23].
@@
For the charmonium system, the leptonic decay constant of $J/\psi$ can be directly determined experimentally [24]. In this work, we calculated the ratios of decay constants $f_{J/\psi}/f_{\eta_{c}}$ and $f^{T}_{J/\psi}/f_{J/\psi}$, where the former can be directly used in the amplitude analysis of $B$ meson decays into charmonium states, and the latter will appear in $b\to c$ decays induced by specific new physics operators. By combining the decay width $\Gamma(J/\psi\to e^{+}e^{-})$ measured experimentally with the ratio $f_{J/\psi}/f_{\eta_{c}}$, we can determine the decay constant $f_{\eta_{c}}$, which is used for the calculation of decay amplitudes $\Gamma(\eta_{c}\to\gamma^{*}\gamma)$ and form factors appearing in $\gamma^{*}\gamma^{*}\to\eta_{c}$ processes [25, 26, 27]. Theoretically, the decay constants $f_{J/\psi}$ and $f_{\eta_{c}}$ were calculated in quenched lattice QCD in [28]. Two-flavor lattice QCD calculations of $f_{J/\psi}$ and $f_{\eta_{c}}$ are given in [29] and [30]. The HPQCD collaboration obtains $f_{J/\psi}$ and/or $f_{\eta_{c}}$ in $2+1$-flavor [31, 32], and $2+1+1$-flavor simulations [33]. The latter work also considers the quenched Quantum Electrodynamics (QED) effects of charm quarks.
@@ 2 Simulation details
The physical mass points of light, strange and charm quarks are fixed by using the meson masses $m_{\pi}^{\rm expt}$, $m_{K}^{\rm expt}$ and $m_{J/\psi}^{\rm expt}$ measured experimentally. However, the corrections from the difference $m_{d}-m_{u}$ and the electromagnetic effects should be removed from the experimental values. The corresponding values are defined to be $m_{\pi}^{\rm phys}\equiv m_{\pi^{0}}^{\rm expt}$ and ($m_{K}^{\rm phys})^{2}\equiv\frac{1}{2}[(m_{K^{+}}^{\rm latt})^{2}+(m_{K^{0}}^{\rm latt})^{2}]$ in isospin-symmetric QCD with $m_{K^{+}}^{\rm latt}=491.405$ MeV and $m_{K^{0}}^{\rm latt}=497.567$ MeV [44, 8, 45]. We then use $m_{\pi}^{\rm phys}=134.98$ MeV, $m_{K}^{\text{phys}}=494.49$ MeV, and $m_{J/\psi}^{\rm phys}=3.0969$ GeV to determine the physical mass points. Note that at the physical pion mass, $D^{*}$ decays into the $P$-wave $D\pi$ state through the strong interaction. In Table 3, we have listed some masses of $D^{(*)}$, with different valence pion masses on our lattice. It is seen that, at each valence pion mass, the mass of the vector meson $D^{*}$ is below the $D\pi$ threshold. Therefore, in this study $D^{*}$ is stable on our lattice. Similarly, $D_{s}^{*}$ is also a stable particle.
@@ 3.1 Meson masses
Figure 2: Left: fitted $aM_{D_{s}}$ as a function of $t_{\text{min}}$ with fixed $t_{\text{max}}=32$. Middle: effective masses (blue points) from the two-point function, compared with $aM_{D_{s}}$ from the fit (orange band) with fitting range $15\leq t\leq 32$. Right: fitted $aM_{D_{s}}$ as a function of the number of states from the multi-state fit in the range $2\leq t\leq 32$.
The right panel of Fig. 2 shows $aM_{D_{s}}$ from the multi-state fits with the fitting range $2\leq t\leq 32$ as a function of the number of states $n$ in the fitting function [46]. After $n$ exceeds $4$, the fitting results stabilize and agree with that from the one-state fit.
@@
[…]
@@
Table 6: Fitted results of $r\equiv m_{\eta_{s}}^{2}/(2m_{K}^{2}-m_{\pi}^{2})$ and $am_{\eta_{s}}$ at various strange quark masses on ensemble f004, with statistical uncertainties estimated by Jackknife analyses.
$am_{s}$ 0.037 0.040 0.043 0.046 0.049 0.052
$r$ 0.996(8) 0.999(8) 1.002(8) 1.005(8) 1.008(7) 1.011(7)
$am_{\eta_{s}}$ 0.2632(7) 0.2739(7) 0.2844(7) 0.2945(7) 0.3043(7) 0.3140(7)
After obtaining the meson masses and decay constants at different valence quark masses on ensemble f004, we used linear interpolation (extrapolation) to obtain the results at the physical mass point. The Particle Data Group [1] and Fermilab Lattice and MILC Collaborations [8] give
$m_{\pi}^{\rm phys}=134.98\ \text{MeV},\quad m_{K}^{\rm phys}=494.49\ \text{MeV},\quad m_{J/\psi}^{\rm phys}=3.0969\ \text{GeV}.$
@@
$(am_{\pi})^{2}_{\rm phys}=0.00321(2)\,,\ (am_{ss})^{2}_{\rm phys}=0.0829(7)\,,\ (am_{J/\psi})_{\rm phys}=1.2996(49)\,.$
@@
$M_{D}^{\text{phys}}=1.866(14)(7)\,\text{GeV},\quad M_{D^{*}}^{\text{phys}}=2.007(24)(8)\,\text{GeV},$ (17)
$M_{D_{s}}^{\text{phys}}=1.9666(46)(75)\,\text{GeV},\quad M_{D_{s}^{*}}^{\text{phys}}=2.1133(71)(80)\,\text{GeV},$
@@ 3.2 Decay constants
$f_{D}=215.3(9.1)(0.8),\quad f_{D^{*}}=223.7(16.2)(0.9),\quad f_{D^{*}}^{T}=190.8(12.6)(1.9)(0.8),$ (21)
$\,f_{D_{s}}=255.7(4.3)(1.0),\quad f_{D_{s}^{*}}=276.8(6.4)(1.1),\quad f_{D_{s}^{*}}^{T}=251.9(6.0)(2.4)(1.0).$
Here, for each of the decay constants, the first error includes statistical and extrapolation/interpolation uncertainties, the last error is from the uncertainty of the lattice spacing, and the error in the middle, if there is one, is due to the uncertainty in the renormalization constant. It is seen in Eq. (8) that the uncertainty of $Z_{V}(=Z_{A})$ can be regarded as negligible. The uncertainty of $Z_{T}$ is around 1% and is much smaller than the first error.
For chiral fermions as used in this work the discretization effects are proportional to the squared lattice spacing $a^{2}$. With three or more lattice spacings we would be able to do linear fittings in $a^{2}$ and obtain results in the continuum limit. The lattice spacing squared corresponding to our current work ($a^{2}\approx 0.007$ fm2) happens to lie almost at the middle point between the continuum limit and the previous work [34] ($a^{2}\approx 0.013$ fm2). A simple linear extrapolation in $a^{2}$ using the results at the two lattice spacings has no degree of freedom and will shift our current result by an amount almost equal to the difference between the two lattice spacings. Therefore, we think the discretization errors can be estimated by simply comparing the results of these two works. For $f_{D_{s}^{(*)}}$, our results are consistent within $1\sigma$ deviation with the previous results , respectively, where $f_{D_{s}}=249(5)\,$MeV and $f_{D_{s}^{*}}=274(5)\,$MeV [34] and we have dropped the discretization error assigned to the previous work. For $f_{D^{(*)}}$, our results are also consistent within $1\sigma$ deviation with the previous results, respectively, where $f_{D}=213(2)\,$MeV and $f_{D^{*}}=234(3)\,$MeV. The differences in the central values are approximately from 1% to $4\%$. Therefore, we take an average value $3\%$ as the estimate of the discretization errors in our decay constants and get (in units of MeV),
$f_{D}=215.3(9.1)(6.5),\quad f_{D^{*}}=223.7(16.3)(6.8),\quad f_{D^{*}}^{T}=190.8(12.8)(5.8),$ (22)
$\,f_{D_{s}}=255.7(4.4)(7.7),\quad f_{D_{s}^{*}}=276.8(6.5)(8.3),\quad f_{D_{s}^{*}}^{T}=251.9(6.6)(7.6),$
@@
$f_{J/\psi}/f_{\eta_{c}}=1.060(15),\quad f^{T}_{J/\psi}/f_{J/\psi}=0.961(10).$ (23)
Here the error includes statistical uncertainty and the one in renormalization constants. Combining these values with the experimental measured $f_{J/\psi}^{\rm expt}=407(4)\,$MeV, which is obtained from its pure leptonic decay width $\Gamma(J/\psi\to e^{+}e^{-})$ [33], we can derive $f_{\eta_{c}}=383.8(6.7)\,$MeV and $f^{T}_{J/\psi}=391.1(5.4)\,$MeV. These results are consistent within $1.5\sigma$ deviation with the results $f_{\eta_{c}}=394.7(2.4)\,$MeV [31], $f^{T}_{J/\psi}=392.7(2.7)\,$MeV and $f^{T}_{J/\psi}/f_{J/\psi}=0.9569(52)$ [51] from the HPQCD Collaboration obtained in the continuum limit. The better agreement in the ratio means the discretization effects are shrunk by the cancellation between the numerator and the denominator. Because we do not have results of decay constants of charmonia from the previous work to compare, we do not give discretization errors for those constants and ratios in this work.